# 3D-GRAND: A Million-Scale Dataset for 3D-LLMs with Better Grounding and Less Hallucination

### v2

@@ 5 Experiments
In this section, we present our experimental setup, including the baselines, datasets, and implementation details. We then report the results of our approach, denoted as 3D-GRAND on the ScanRefer [10] and the 3D-POPE benchmark, demonstrating the effectiveness in improving grounding and reducing hallucination. Finally, we conduct an ablation study to analyze the impact of different components of our model and training strategy.
@@ 5.1 Experimental Setup
Baselines. We compare our 3D-GRAND against the following baselines: 3D-LLM [27], LEO [30], and 3D-Vista [83]. Each model, along with the specific checkpoint used to obtain the results, is documented in the appendix. Our proposed model is based on Llama-2 [64]. The input is object-centric context, including a scene graph with each object’s category, centroid (x, y, z), and extent (width, height, depth), along with the text instruction and user query. During training, we utilized ground-truth centroids and extents. For inference, we used bounding boxes predicted by Mask3D [59]. Examples of input/output and details of the model can be found in the supplementary material.
Datasets. We evaluate our model 3D-GRAND on two datasets: 3D-POPE and ScanRefer. 3D-POPE is our newly introduced benchmark dataset for evaluating object hallucination in 3D-LLMs, as described in Section 4. For ScanRefer, We utilized the validation split which contains 9,508 natural language descriptions of 2,068 objects in 141 ScanNet [14] scenes.
@@
Implementation Details. The 3D-GRAND model is LoRA-finetuned [28] based off Llama-2. We use DeepSpeed ZeRO-2 [55] and FlashAttention [16] to save GPU memory and speed up training. The model is trained in BF16 precision on 12 NVIDIA A40 GPUs with a combined batch size of 96 and a learning rate of 2e-4. We use the AdamW [45] optimizer with a weight decay of 0.01 and a cosine learning rate scheduler. We train the mode for 10k steps, which takes approximately 48 hours.
@@ 5.2 Results on 3D-POPE
We first evaluate these approaches on 3D-POPE and report results on Table 2. Results show that 3D-LLM [27] almost always produces yes responses to any question. 3D-VisTA [83] performs similarly to the random baseline. LEO [30] tends to answer yes frequently, but its precision indicates a similar object hallucination rate to the random baseline. In our evaluation, 3D-GRAND achieved exceptional performance, with 93.34% precision and 89.12% accuracy when tested with random sampling. However, our model struggles with the more difficult splits, Popular and Adversarial, which demonstrates the effectiveness and rigorousness of 3D-POPE as a benchmark. Moreover, we emphasize that our model has never encountered ScanNet during training. More analysis on 3D hallucination can be found in the supplementary material.
@@ 5.3 Results on ScanRefer
We report results on ScanRefer in Table 4. Our model significantly outperforms all the baselines, achieving state-of-the-art zero-shot results on both Acc@0.25 and Acc@0.5. Notably, our model surpasses the previous best-performing model, 3D-LLM, by 7.7% on accuracy@0.25IoU. We emphasize that our model, unlike 3D-LLM, has never seen ScanNet scenes during its training (zero-shot) and is only trained on synthetics 3D scene instead of real scans. Therefore, this results provide a promising early signal that sim-to-real transfer can be achieved via our dense-grounded large-scale dataset.
@@
(a) Grounding capability. Higher is better.
(b) Hallucination rate. Lower is better.
Figure 4: 3D-GRAND data scaling analysis on zero-shot, sim-to-real grounding capability, and hallucination. Grounding performance consistently improves as data scales up. Model trained with densely-grounded data exhibits better grounding capability compared to that trained without. Additionally, model hallucinates less when exposed to more data from 3D-GRAND. Here, the Hallucination Rate is calculated as $(1-\text{Precision})$ on 3D-POPE.
@@ 5.4 Ablation Study
To better understand the impact of different components of our 3D-LLM, we conduct an ablation study on the ScanRefer and 3D-POPE benchmarks.
Grounding tokens. We show the results of our model with different types of grounding methods in Table 5. We also show results on 3D-POPE in Table 4. In general, the model has a worse grounding performance and more hallucinations without grounding tokens. “Ground First” and “ground later” refer to whether the dense grounding (grounding every single object mentioned) of the object reference query happens before or after the model outputs the final answer for the refer expression. The former effectively constitutes a chain-of-thought reasoning process [67], which is likely why the performance increases compared to the latter. See Appendix for details.
Mask3D proposals. Finally, we show the upper bound of our approach in Table 5. Our results are based on Mask3D proposals. Due to the context length of LLM, we only use top-40 proposals.
@@ 5.5 Data Scaling and Sim-to-Real Transfer
The results are presented in Figure 4. Our model is trained on synthetic 3D scenes from 3D-FRONT and Structured3D [79, 23], and evaluated on real-world 3D scans from ScanNet [14]. The grounding performance consistently improves, and the hallucination rate drops as the densely-grounded data scales up. Notably, our model trained on densely grounded data scales better than the same model trained without such data. These findings pave the way for a future where we can scale 3D-text understanding using synthetic scenes obtained from simulation, which is much cheaper and more accessible to obtain.
@@ 6 Conclusion
In this paper, we introduced 3D-GRAND, a large-scale, densely-grounded 3D-text dataset designed for grounded 3D instruction tuning, and 3D-POPE, a comprehensive benchmark for evaluating object hallucination in 3D-LLMs. Through extensive experiments, we demonstrated the effectiveness of our dataset on 3D-LLMs in improving grounding and reducing hallucination, achieving state-of-the-art performance on the ScanRefer and 3D-POPE benchmarks. Our ablation study and qualitative analysis highlighted the importance of densely-grounded instruction tuning, the data scaling law, and effective sim-to-real transfer in developing high-performing 3D-LLMs. We hope our contributions and findings can spark further research and innovation in this field, ultimately leading to the development of more advanced and capable 3D-LLMs for a wide range of applications.
@@ 7 Acknowledgement
Appendix A Model Details
A.1 Model input and output demonstration
@@
A.2 Training Data
There are two flavors of models that we fine-tuned: one grounded object reference model, and one grounded QA model. The grounded object reference model was trained using the grounded object reference data on 3D-FRONT train split, which consist of 234,791 3D-text pairs, each of which are densely grounded. This model was used to generate the ScanRefer results presented in Table 4, 5, and Figure 4 The grounded QA model was trained using a subset of 200k grounded QA: object existence data from the 3D-FRONT train split. The reason that we select a subset of 200k QAs is simply because the entire grounded QA dataset is too large and we do not have enough resource to train on all data. However, as shown in Table 2 and Figure 4, we find even such a subset is very effective in reducing hallucinations in 3D-LLMs.
@@ A.3 Training Details
Appendix B Additional 3D-GRAND Data Collection
B.1 Point Cloud Generation Pipeline for 3D-Front
@@
Appendix C Addtional 3D POPE results
C.1 3D Pope Results on NYU40
@@ Appendix C Addtional 3D POPE results
Appendix D Human Validation
@@ D.1.1 Task Generation
D.1.2 Crowd-sourcing Validity
@@
Before presenting any tasks to the workers, we present them with a set of instruction tasks that show an example annotation, the correct response (as determined by us), and the reason why that response is correct. They are paired with an incorrect example and an explanation of why it is incorrect in order to ensure unbiased annotations. Examples of qualifying instructions are shown in 11. These instructions are intentionally brief, as we found through trial-and-error that longer, paragraph-based instructions were largely ignored by annotators.

### v3

@@ 5 Experiments
We present our experimental setup, including baselines, datasets, and implementation details. We then report the results of our approach, denoted as 3D-GRAND on ScanRefer [10] and the 3D-POPE benchmarks, demonstrating the effectiveness in improving grounding and reducing hallucination. Finally, we report an ablation study analyzing the impact of components of our model and training strategy.
@@ 5.1 Experimental Setup
Model. Our proposed model is based on Llama-2 [66]. The input is object-centric context, including a scene graph with each object’s category, centroid (x, y, z), and extent (width, height, depth), along with the text instruction and user query. During training, we utilized ground-truth centroids and extents. For inference, we used bounding boxes predicted by Mask3D [61]. Examples of input/output and details of the model can be found in the supplementary material.
Baselines. We compare our 3D-GRAND against the following baselines: 3D-LLM [28], LEO [31], and 3D-Vista [85]. Each model, along with the specific checkpoint used to obtain the results, is documented in the appendix.
Datasets. We evaluate our model 3D-GRAND on 3D-POPE and ScanRefer. 3D-POPE is our newly introduced benchmark dataset for evaluating object hallucination in 3D-LLMs, as described in Section 3. For ScanRefer, We utilized the validation split which contains 9,508 natural language descriptions of 2,068 objects in 141 ScanNet [15] scenes.
@@
[…]
@@ 5.2 Results on 3D-POPE
We first evaluate these approaches on 3D-POPE, with results in Table 3. Results show that 3D-LLM [28] almost always produces yes responses to any question. 3D-VisTA [85] performs similarly to the random baseline. LEO [31] tends to answer yes frequently, but its precision indicates a similar object hallucination rate to the random baseline. In our evaluation, 3D-GRAND achieved better performance, with 93.34% precision and 89.12% accuracy when tested with random sampling. Our model struggles with the more difficult splits, Popular and Adversarial, which demonstrates the effectiveness and rigorousness of 3D-POPE as a benchmark. Moreover, we emphasize that our model has never encountered ScanNet during training. More analysis on 3D hallucination can be found in the supplementary material.
Model Generative 3D-LLM? Never Seen ScanNet? Acc@0.25 Acc@0.5
Non-LLM based
ScanRefer 37.3 24.3
MVT 40.8 33.3
3DVG-Trans 45.9 34.5
ViL3DRel 47.9 37.7
M3DRef-CLIP 51.9 44.7
Non-Generative 3D-LLMs
3D-VisTA (zero-shot) 33.2 29.6
SceneVerse (zero-shot) 35.2 31.1
Generative 3D-LLMs
3D-LLM 30.3 -
LLM-Grounder 17.1 5.3
3D-GRAND (Ours) 38.0 27.4
Table 5: ScanRefer Results for visual grounding capability of 3D-LLMs. 3D-GRAND achieves the best zero-shot performance among 3D-LLMs, providing signals for sim-to-real transfer.
@@ 5.3 Results on ScanRefer
We report results on ScanRefer in Table 5. There are a few important observations on this result:
Our 3D-LLM trained with 3D-GRAND data achieved the best Acc@0.25 among all models. Notably, our model surpasses the previous best-performing model, 3D-LLM, by 7.7% on accuracy@0.25IoU. We emphasize that our model, unlike 3D-LLM, has never seen ScanNet scenes during its training (zero-shot) and is only trained on synthetic 3D scenes instead of real scans. Therefore, these results provide a promising early signal that sim-to-real transfer can be achieved via our densely-grounded large-scale dataset.
Our generative 3D-LLM model (one that a user can chat with) performs better or on par compared to non-generative 3D-LLMs such as 3D-VisTA and SceneVerse. In the past, generative 3D-LLMs are usually significantly outperformed by non-generative 3D-LLMs, as the latter usually sacrificed the ability to chat in exchange for incorporating specialized model designs, such as producing scores for each object candidate. These designs are closer to traditional non-LLM-based specialized models. But here, we observe that the gap between the two modeling choices is closing with the help of large-scale densely-grounded data like 3D-GRAND.
Our model is just a deliberately a primarily text-based model (Sec. 5.1) to demonstrate the effectiveness of the dataset - little visual information is conveyed in our model between the mask proposal to the LLM, in contrast to other sophisticated models where 3D object embeddings are used to better represent visual information. Thus 3D-GRAND has more potential to be unlocked in the future.
@@
Figure 3: Data scaling analysis on zero-shot, sim-to-real grounding capability, and hallucination. Grounding performance (left two subfigures) consistently improves as data scales up. Model trained with densely-grounded data exhibits better grounding capability compared to that trained without. Additionally (right subfigure), the model hallucinates less when exposed to more data from 3D-GRAND. Here, the Hallucination Rate is calculated as $(1-\text{Precision})$ on 3D-POPE.
@@ 5.4 Ablation Study
To better understand the components of our 3D-LLM, we conduct an ablation study on ScanRefer and 3D-POPE.
3D-POPE Model Precision
Random 3D-GRAND 93.34
w/o grounding tokens (-1.38)
Popular 3D-GRAND 73.05
w/o grounding tokens (-2.68)
Adversarial 3D-GRAND 69.86
w/o grounding tokens (-2.38)
Grounding tokens. Table 6 shows results of our model with different types of grounding methods. We also show results on 3D-POPE in Table 7. In general, the model has a worse grounding performance and more hallucinations without grounding tokens. “Ground First” and “Ground Later” refer to whether the dense grounding (grounding every single object mentioned) of the object reference query happens before or after the model outputs the final answer for the refer expression. The former functions like a chain-of-thought reasoning process [69], which is likely why the performance increases compared to the latter. See Appendix for details.
Mask3D proposals. Finally, we show the upper bound of our approach in Table 6, with Mask3D proposals. Due to the LLM context length, we only use top-40 proposals.
@@ 5.5 Data Scaling and Sim-to-Real Transfer
We present results in Figure 3. Our model is trained on synthetic 3D scenes from 3D-FRONT and Structured3D [81, 24], and evaluated on real-world 3D scans from ScanNet [15]. Grounding performance consistently improves and the hallucination rate drops as the densely-grounded data scales up. Notably, our model trained on densely grounded data scales better than the same model trained without such data. These findings pave the way for scaling 3D-text understanding using synthetic scenes obtained from simulation, which is cheaper and easier to obtain.
@@ 6 Conclusion
We introduced 3D-GRAND, a large-scale, densely-grounded 3D-text dataset designed for grounded 3D instruction tuning, and 3D-POPE, a comprehensive benchmark for evaluating object hallucination in 3D-LLMs. Through extensive experiments, we demonstrated the effectiveness of our dataset on 3D-LLMs in improving grounding and reducing hallucination, achieving state-of-the-art performance on the ScanRefer and 3D-POPE benchmarks. Our ablation analysis demonstrated the importance of densely-grounded instruction tuning, data scaling laws, and effective sim-to-real transfer in developing high-performing 3D-LLMs. We hope our findings can spark further research and innovation in this field, leading to the development of more advanced and capable 3D-LLMs for a wide range of applications.
@@ 7 Acknowledgement
Supplementary Material
Appendix A Detailed Comparison with SceneVerse
Among recent datasets, 3D-GRAND is most similar to SceneVerse [34]. They are both million-scale grounding datasets for 3D-LLMs. However, there are a few key differences: (1) SceneVerse [34] provides only sparse grounding, while 3D-GRAND is densely grounded. In 3D-GRAND, every noun phrase in the text—whether it’s in captions, QAs, or object references—is explicitly grounded to a corresponding object in the 3D scene, whereas SceneVerse does not offer this level of grounding granularity. To elucidate this difference, we present Table 2 and 8 that compares SceneVerse and 3D-GRAND; (2) the language annotations of 3D-GRANDare more trustworthy and have higher quality. Hallucination is known as one of the most common mistakes of LLMs [32, 43, 58] In 3D-GRAND, we employ a hallucination filter to check and delete any annotations with hallucinated object IDs. This is not possible for SceneVerse since they have pure language output. 3D-GRAND is also quality-checked by humans to ensure the quality.
Grounding Granularity Object Reference Data
SceneVerse Session-level many-to-one grounding
This is a big cotton sofa. It is between the window and the
wooden table. $\rightarrow$ sofa-3
3D-GRAND Noun-level one-to-one grounding
This is a big cotton sofa [sofa-3]. It [sofa-3] is
between the window [window-0] and wooden table [table-4].
Table 8: Example of grounding granularity. 3D-GRAND focuses on dense grounding.
Definitions of grounding granularity:
  • Paragraph-level set-to-set grounding: Many sentences in a long paragraph, each containing several object nouns, are linked to a set of 3D objects without clear associations from specific sentences/noun phrases to objects.
  • Session-level many-to-one grounding: Multiple sentences in one session, where each sentence can describe several objects (targets and landmarks), are associated with one 3D object.
  • Noun-level one-to-one grounding: Each noun phrase in each sentence is explicitly matched with one 3D object.
Appendix B Implementation of Our Model
B.1 Model input and output demonstration
@@
B.2 Training Data
There are two flavors of models that we fine-tuned: one grounded object reference model, and one grounded QA model. The grounded object reference model was trained using the grounded object reference data on 3D-FRONT train split, which consist of 234,791 3D-text pairs, each of which are densely grounded. This model was used to generate the ScanRefer results presented in Table 5, 6, and Figure 3 The grounded QA model was trained using a subset of 200k grounded QA: object existence data from the 3D-FRONT train split. The reason that we select a subset of 200k QAs is simply because the entire grounded QA dataset is too large and we do not have enough resource to train on all data. However, as shown in Table 3 and Figure 3, we find even such a subset is very effective in reducing hallucinations in 3D-LLMs.
@@ B.3 Training Details
Appendix C Additional 3D-GRAND Data Collection
C.1 Point Cloud Generation Pipeline for 3D-Front
@@
Appendix D Additional 3D POPE results
D.1 3D Pope Results on NYU40
@@ Appendix D Additional 3D POPE results
Appendix E Human Validation
@@ E.1.1 Task Generation
E.1.2 Crowd-sourcing Validity
@@
Before presenting any tasks to the workers, we present them with a set of instruction tasks that show an example annotation, the correct response (as determined by us), and the reason why that response is correct. They are paired with an incorrect example and an explanation of why it is incorrect in order to ensure unbiased annotations. Examples of qualifying instructions are shown in 10. These instructions are intentionally brief, as we found through trial-and-error that longer, paragraph-based instructions were largely ignored by annotators.